Determine the complete power information of a series circuit composed of R=17.32 Ω and having pf=0.866 lag, for an applied voltage v(t)=100*sqrt(2)*cos(1000t + 30)
Answer #1

ceries RL - cuit * you know, Generally anductor can absorb the Logging Reactive power series RL- circuit alts as logging ReciIn this problem Here that power factor =colo = 0.866 lagging the power factor is lagging, it means. the circuit is series RL-1000xL 17.32 17.32 x 103 13 [2 = 10mt through direit Cevyen vel. I = 1002 130 RAIN2 Rtite I - 1000 L30 17.32.7110 I = 7.07120S = -x look (30 87.07120)* s = f x 10050 130 X 7.07120 s = 433.01 + j2u9.99.7 S = 433.01 +1250 s = ptig L Realtive power deliyou know already Resistor can deliver the artive power * R.m-s value of current for cosine tunition IRMS = I maximum - = (707